Suppose that X ∼ N(-1.4,3.1), Y ∼ N(2.9,1.7), and Z ∼ N(1.1, 0.4) are independent random variables.

Find the probability that 4.8X + 3Y + 4Z ≥ 9.6.

Round your answer to the nearest thousandth.
Answer #1

Ans:

E(4.8X + 3Y + 4Z)=-1.4*4.8+3*2.9+4*1.1=6.38

SD(4.8X + 3Y + 4Z)=SQRT(4.8^2*3.1+3^2*1.7+4^2*0.4)=9.65

z=(9.6-6.38)/9.65

z=0.334

P(z>=0.334)=0.369